Daily Events on 2008/09/16 (Tue)

last updated at  2008/09/16 14:23:26 JST


<< prev day

@

next day >>


Station

Pass No.
20080916-

Contact time (UT)

Max. Elev.
 (deg)

Events

AOS

LOS

MSP1

8000

06:46:06

06:59:32

47.3

AOCS: Upload pointing table
OP/OG upload
XRT: TI-CMD
EIS: TI-CMD
SOT: TI-CMD

USC20

0100

07:52:48

08:05:13

22.7

OP/OG upload, XRT:TI-CMD, SOT:TI-CMD, EIS:TI-CMD (backup)
AOCS: Upload orbital element
DR playback (by OP)
EIS: Table Upload

USC20

0200

09:29:46

09:41:59

26.9

DR playback (by OP)
XRT: Table upload
SOT: Table upload

USC34

0400

21:31:14

21:44:48

74.4

DR playback (by OP)

Flare list(LM) : http://www.lmsal.com/solarsoft/last_events/


DR information

Remaining amount
(prediction)

- Mbits @ OP_START ( )


SOT

XRT

EIS

DR allocation [%]

70

15

15

DR allocation [Mbits]

-

-

-

DR TLMLMT [Mbits]

-

-

-


daily note

SOT requests AOCS pointing table upload in first pass, to change pointing for VTT QS observations before the OP start time

XRT:
- Support HOP73 - Filament.
- Support VTT QS observations on Tenerife

SOT:
- Support HOP 73: Quiescent Prominence or Filament
- Quiet Sun coordination with VTT

EIS:
- Grating focus adjustment will be done this week. Requires ~14 hours of North Pole pointing for three consecutive days starting after the end of HOP73 at 12UT. The plan is to do this starting from Tuesday upload.
- Support HOP73



CHIEF
TOHBAN

SUB
TOHBAN

CP

SOT CO

XRT CO

EIS CO

EVENING

Sugita

Magara

Tarbell

Matsumoto

David H Brooks

MORNING


operation files

pointing file : none